Grade 12th passMechanics

We use P=FV only when velocity is constant? When a body accelerates then Power=Fa×d/t

Hari Krishna Nethi
9 Years agoGrade 12th pass

1 Answer

somi teez
9 Years ago
hi,
to some extent you are right..but not exactly see
P=F*V 
IF you right in terms of acceleration you can write P=F*V=M*A*V=p*A
here, 
p=momentum,A=acceleration, M=mass